# Reduced Rank Regression for Mixed Predictor and Response Variables

### v1

@@ 6 Conclusions and Discussion
Furthermore, we also considered different types of predictor variables, whereas the earlier models could only deal with numeric variables. Therefore, discrete predictor variables have to be coded as dummy variables. We, however, used optimal scaling to quantify the discrete predictors. The quantifications are chosen such that they maximize the log likelihood of the final model.
We developed a majorization-minimization algorithm for maximum likelihood estimation of the parameters of the model. We showed that the negative log likelihood for the different response variables can be majorized by a least squares function. For minimizing least squares functions a lot of knowledge is available (cf. Ten Berge,, 1993). The algorithm monotonically converges to a minimum. Because each of the original functions per response variable is convex, the sum of these is also convex. Therefore, the attained minimum is also the global minimum. In our application, the algorithm turned out to be quite fast, that is, only a few iterations are needed for convergence. Other properties of the algorithm still have to be investigated.
@@
Other future possibilities include the extension of the model to high dimensional predictors (i.e., large $P$) and high dimensional responses (i.e., large $R$). In that case, penalties such as the lasso (Tibshirani,, 1996), ridge (Hoerl and Kennard,, 1970), or the group lasso (Yuan and Lin,, 2006) to penalize the regression weights and/or the loadings. Of course, this requires the development and testing of new algorithms and ways to find optimal penalty parameters.
@@
A final future extension, is the generalization of our GMR3 method to nested or longitudinal data. We assumed the observations to be independent, but in many applications certain observations might be correlated, for example when observe children in schools, or when we follow participants over time. We need to take the dependencies into account, which can be done by including for example random effects as in multilevel models (Snijders and Bosker,, 2011; Kreft and De Leeuw,, 1998), or by adapting the information criteria (Pan,, 2001) and the bootstrap (Sherman and Cessie,, 1997; Deen and de Rooij,, 2020). Often, in longitudinal data analysis the interest lies in the development over time for different groups. How to answer such a question when we include optimal scaling of the group variable is something that needs to be investigated.
Statements and Declarations
Data availability:
The Eurobarometer data are freely available from the the website of the European Union https://europa.eu/eurobarometer/screen/home. R-code for the analyses of these data sets is available on the github-page of the first author.
Funding:
The first author (MdR) is part of the GUTS program which is funded by an NWO Gravitation programme supported by the Dutch Ministry of Education, Culture and Science of the government of the Netherlands, Grant nr 024.005.011. The second and third author (LC and RS) are funded by FIT4MEDROB PNC0000007.
Competing Interests:
The authors have no relevant financial or non-financial interests to disclose.
Author Contributions:
MdR developed the original methodology described in this paper and the MM algorithm, wrote main parts of the text. LC retrieved the data and performed all data analyses. RS wrote the section about model selection. All authors reviewed the first version of the paper and contributed to the improvement of the final version.

### v2

@@ 7 Conclusions and Discussion
Furthermore, we also considered different types of predictor variables, whereas the earlier models could only deal with numeric variables. Therefore, discrete predictor variables have to be coded as dummy variables. We, however, used optimal scaling to quantify the discrete predictors. Whereas there is a lot of knowledge on optimal scaling within least squares problems (see, for example, Gifi,, 1990), optimal scaling has not been applied within maximum likelihood estimation. We show an estimation procedure, where the quantifications are chosen such that they maximize the log likelihood of the final model. This was made possible by the employment of the MM algorithm we developed.
We developed a majorization-minimization algorithm for maximum likelihood estimation of the parameters of the model. We showed that the negative log likelihood for the different response variables can be majorized by a least squares function. For minimizing least squares functions a lot of knowledge is available (cf. Ten Berge,, 1993). The algorithm monotonically converges to a minimum. Because each of the original functions per response variable is convex, the sum of these is also convex. Therefore, the attained minimum is also the global minimum. In our application, the algorithm turned out to be quite fast, that is, only a few iterations are needed for convergence. For our cross validation and bootstrap functions we used warm starts to speed up the procedures, where the starting values are equal to the final solutions obtained on the complete data set. We discussed in Section 3.1 that the algorithm might have difficulties in finding the minimum when the estimated variance of the residuals for the numeric response variables is small. In our application and simulation studies we did not encounter this issue. Other properties of the algorithm still have to be investigated.
In Section 5.2, we proposed a step-wise model selection procedure where first the rank or dimensionality is chosen and thereafter a bootstrap procedure is employed to verify which predictor variables have an effect on the responses. Such a step-wise procedure is efficient in terms of the number of models to be fitted. However, it does not guarantee to find the best model. Therefore, it would be better to fit models wit varying sets of predictors in all possible dimensionalities. Such a procedure requires fitting many models and can be time consuming. Still, if we would use the full grid of possible models, different statistics of model selection (AIC and BIC, for example) would probably point towards different best models. In this context, it is also good to point out that we generally do not believe in a true underlying model that we need to recover. Instead, we search for a model that can describes the data relatively good and is stable.
We applied our method to an empirical data set and compared the results to the results obtained using standard separate regression models for each of the responses. We showed that these standard regression models do not take into account the ordered nature of the predictor variables (as they use dummy variables for the categorical predictors), whereas our model does. When fitting separate models, we need many more parameters than in our reduced rank model and we showed that this leads to overfitting, that is, large and unstable parameter estimates. In contrast, our GMR3 model leads to stable results. Furthermore, our model clearly show which response variables are affected in a similar way by the predictors and our model can take into account the ordered nature of some of the predictors.
@@
Other future possibilities include the extension of the model to high dimensional predictors (i.e., large $P$) and high dimensional responses (i.e., large $R$). In that case, penalties such as the lasso (Tibshirani,, 1996), ridge (Hoerl and Kennard,, 1970), or the group lasso (Yuan and Lin,, 2006) to penalize the regression weights and/or the loadings. Of course, this requires the development and testing of new algorithms and ways to find optimal penalty parameters. A final future extension, is the generalization of our GMR3 method to nested or longitudinal data. We assumed the observations to be independent, but in many applications certain observations might be correlated, for example when observe children in schools, or when we follow participants over time. We need to take the dependencies into account, which can be done by including for example random effects as in multilevel models (Snijders and Bosker,, 2011; Kreft and De Leeuw,, 1998), or by adapting the information criteria (Pan,, 2001) and the bootstrap (Sherman and Cessie,, 1997; Deen and de Rooij,, 2020). Often, in longitudinal data analysis the interest lies in the development over time for different groups. How to answer such a question when we include optimal scaling of the group variable is something that needs to be investigated.
Appendix A: comparison to separate models
We fitted standard binary and ordinal logistic regression models to the seven response variables. As predictors, we used age, political alignment, gender, degree of urbanization, and education level. The categorical predictor variables were coded using dummies, with the first category as baseline. As, all predictors were of importance in our gmr3 model, we did not do any further model selection. The sum of deviances of these seven models is 9923.01 (corresponding to $\mathcal{L}$ of 4961.51). Of course this value is lower compared to our analysis, because more parameters are fitted. That is, the total number of parameters is 115. The corresponding AIC equals 10153.01, while the BIC is 10696.94. Both values are higher than those for our selected model.
The parameter estimates obtained using the seven separate models are given in Table 2.
T FE CI MW FS DI RE
A -0.18 -0.32 -0.04 0.09 0.16 0.46 0.25
PA2 -0.96 -0.36 -0.81 -0.59 -0.56 0.17 -0.59
PA3 -1.30 -0.98 -1.40 -0.99 -0.72 0.35 -1.04
G2 -0.18 0.02 -0.22 0.02 -0.40 -0.53 -0.21
U2 0.06 -0.01 0.03 -0.04 0.11 -0.00 0.02
U3 0.69 0.15 0.37 0.18 0.40 -0.07 0.41
E2 -0.71 0.12 -1.54 -10.14 -13.26 -12.48 -11.87
E3 -0.35 0.33 -0.99 -11.19 -12.36 -12.99 -11.88
E4 -0.26 0.04 -0.98 -10.92 -12.03 -13.22 -12.14
E5 -0.13 2.24 -0.05 -11.02 -12.05 -12.51 -11.27
E6 0.22 -0.07 -0.71 -11.08 -12.21 -13.19 -11.85
E7 0.09 0.14 -0.21 -11.19 -12.00 -13.46 -11.93
E8 0.65 0.14 -0.18 -11.43 -11.71 -13.18 -12.04
E9 1.52 0.87 0.51 -10.14 -10.04 -13.02 -12.55
Table 2: Coefficients of separately fitted logistic or ordinal logistic regression models
To compare the fitted parameters of the separate models to those of our model, we need to compute some implied coefficients. The regression coefficients corresponding to the dummy variables, specify a difference in the log odds between two categories of a predictor variable. We can obtain similar coefficients from our model. We already saw the implied coefficients $\bm{BV}^{\prime}$. To obtain the effect of two predictor categories, we need to incorporate the optimal scaling. Suppose, we like to know the effect of the second category against the first of the first predictor on the different response variables. Then we need to compute $(\phi_{1}(2)-\phi_{1}(1))\bm{b}^{\prime}_{1}\bm{V}^{\prime}$. Here $\bm{b}_{1}$ is the $S$-dimensional vector of coefficients for the first predictor, and $\phi_{1}(1)$ is the optimally scaled value for category 1 of the first predictor variable and $\phi_{1}(2)$ the optimally scaled value for category 2 of the first predictor variable. Similarly, for the third against the first category of the first predictor we need $(\phi_{1}(3)-\phi_{1}(1))\bm{b}^{\prime}_{1}\bm{V}^{\prime}$. In this way, we can compare the effects obtained in the separate models to the effects obtained in our model. The implied coefficients are given in Table 3
T FE CI MW FS DI RE
A -0.16 -0.27 -0.05 0.06 0.17 0.50 0.23
PA2 -0.96 -0.42 -0.90 -0.33 -0.57 0.16 -0.46
PA3 -1.55 -0.68 -1.45 -0.53 -0.92 0.26 -0.75
G2 -0.12 0.13 -0.21 -0.14 -0.31 -0.41 -0.34
U2 0.04 0.02 0.04 0.02 0.03 -0.00 0.02
U3 0.53 0.20 0.51 0.20 0.35 -0.02 0.30
E2 0.00 0.00 0.00 0.00 0.00 -0.00 0.00
E3 0.28 0.13 0.25 0.09 0.15 -0.08 0.11
E4 0.33 0.16 0.30 0.11 0.18 -0.09 0.14
E5 0.64 0.31 0.58 0.20 0.34 -0.18 0.26
E6 0.64 0.31 0.58 0.20 0.34 -0.18 0.26
E7 0.81 0.39 0.74 0.26 0.43 -0.23 0.33
E8 0.97 0.47 0.89 0.31 0.51 -0.27 0.40
E9 1.77 0.86 1.62 0.56 0.94 -0.49 0.73
Table 3: Implied coefficients of GMR3 model
Whereas some of the coefficients of the separate models are really large, we see that all our coefficients behave well. Because, we assumed some predictor variables to be ordinal, we see that the implied coefficients are neatly ordered, whereas such ordering is absent from the fitted separate models.
We can also look at the standard errors of these coefficients. Applying the bootstrap, and computing these coefficients in each of the bootstrap samples, we can obtain a bootstrap standard error by computing the standard deviation over the bootstrap samples. For the separately fitted models these standard errors are given in Table 4. The corresponding standard errors of our model are given in Table 5. It can be verified that the results of our model are much more stable, that is, the standard errors are much smaller overall. In other words, the efficiency of the our estimator is much higher, specifically for the categorical predictors.
T FE CI MW FS DI RE
A 0.08 0.08 0.07 0.08 0.08 0.07 0.07
PA2 0.20 0.18 0.16 0.19 0.20 0.16 0.16
PA3 0.22 0.21 0.21 0.21 0.24 0.19 0.20
G2 0.17 0.16 0.15 0.17 0.18 0.15 0.15
U2 0.18 0.17 0.16 0.18 0.19 0.16 0.17
U3 0.22 0.20 0.19 0.21 0.24 0.20 0.21
E2 11.17 10.79 0.84 4.76 3.62 4.61 4.47
E3 10.08 9.96 0.42 3.22 3.55 4.27 4.37
E4 10.09 9.97 0.40 3.22 3.55 4.27 4.36
E5 10.86 11.88 0.87 5.25 6.18 6.16 6.35
E6 10.10 9.99 0.45 3.24 3.58 4.29 4.36
E7 10.09 9.98 0.40 3.23 3.55 4.28 4.38
E8 10.10 9.98 0.40 3.22 3.55 4.27 4.36
E9 10.71 10.08 0.52 2.56 5.34 0.86 0.99
Table 4: Bootstrap standard error estimates of separately fitted models
T FE CI MW FS DI RE
A 0.09 0.08 0.07 0.10 0.08 0.08 0.09
PA2 0.16 0.10 0.13 0.13 0.13 0.12 0.14
PA3 0.20 0.17 0.19 0.20 0.17 0.20 0.21
G2 0.14 0.11 0.13 0.10 0.14 0.15 0.11
U2 0.10 0.04 0.09 0.04 0.07 0.04 0.06
U3 0.17 0.11 0.15 0.11 0.15 0.16 0.15
E2 0.37 0.21 0.31 0.09 0.16 0.22 0.12
E3 0.52 0.30 0.43 0.13 0.24 0.32 0.17
E4 0.54 0.32 0.45 0.14 0.26 0.35 0.20
E5 0.54 0.33 0.45 0.15 0.26 0.38 0.21
E6 0.54 0.33 0.45 0.16 0.27 0.39 0.22
E7 0.54 0.35 0.45 0.17 0.28 0.44 0.24
E8 0.64 0.39 0.53 0.22 0.37 0.50 0.31
E9 0.70 0.48 0.57 0.31 0.46 0.64 0.40
Table 5: Bootstrap standard error estimates of GMR3 models